# First Author Affiliation / Address line 1 Affiliation / Address line 2

### v2

@@ Abstract
Ziwen Xu1, Shuxun Wang1, Kewei Xu1, Haoming Xu1, Mengru Wang1, Xinle Deng1, Yunzhi Yao1, Guozhou Zheng1, Huajun Chen1, Ningyu Zhang1thanks:  Corresponding author. 1Zhejiang University {ziwen.xu,zhangningyu}@zju.edu.cn [Uncaptioned image]https://zjunlp.github.io/project/EasyEdit2
@@ 3.7 Evaluators Module
Language Feature Which club is Messi at? Lionel Messi currently plays for **Inter Miami CF** in Major League Soccer (MLS). {CJK} UTF8gbsn梅西目前效力于 **迈阿密国际足球俱乐部** (Inter MiamiCF)。
@@ 6 Conclusion and Future Work
Steering techniques significantly influence model behavior during the inference process. When this can be beneficial, deliberately steering in a negative direction risks generating unethical or harmful content, violating fundamental ethical principles and raising concerns about appropriate application. To avoid these risks, rigorous safety inspections and ethical safeguards must be prioritized when using EasyEdit2 to steer model.
@@ Acknowledgements
Our sincerest thanks are extended to CAA [^6] https://github.com/nrimsky/CAA, LM-Steer [^7] https://github.com/Glaciohound/LM-Steer, and AxBench [^8] https://github.com/stanfordnlp/axbench for their invaluable contributions to our project. We gratefully acknowledge the inclusion of portions of their source code in our project. We also extend our sincere thanks to the community for its ongoing support and collaboration. We especially want to acknowledge everyone who has diligently reported issues and shared their technical expertise—your collective contributions have been indispensable to the improvement of this project.
@@ B.2 Adjustability of Steering Vectors
We investigate the adjustability of merged steering vectors on the Gemma-2-9b model by modulating their effect with different multipliers during inference. Using CAA-based vectors from safety and sentiment tasks, we test three merging strategies: Linear, TIES, and DARE-TIES. Details of these merging strategies are provided in Section 3.4. Each steering vector is scaled by a multiplier from $-2$ to $2$ and applied during inference.

### v3

@@ Abstract
Ziwen Xu1, Shuxun Wang1, Kewei Xu1, Haoming Xu1, Mengru Wang1, Xinle Deng1, Yunzhi Yao1, Guozhou Zheng2, Huajun Chen1, Ningyu Zhang1thanks:  Corresponding author. 1Zhejiang University 2Ocean Research Center of Zhoushan, Zhejiang University {ziwen.xu,zhangningyu}@zju.edu.cn [Uncaptioned image]https://zjunlp.github.io/project/EasyEdit2
@@ 3.7 Evaluators Module
Language Feature Which club is Messi at? Lionel Messi currently plays for **Inter Miami CF** in Major League Soccer (MLS). 梅西目前效力于 **迈阿密国际足球俱乐部** (Inter MiamiCF)。
@@ 6 Conclusion and Future Work
Steering techniques can beneficially adjust LLM behavior, but also pose risks of misuse. As shown in Table 1, inappropriate steering may degrade safety, and malicious use could deliberately induce unethical or harmful content. To mitigate such risks, EasyEdit2 should be applied with curated steering data, systematic safety evaluation, and restricted access to harmful configurations. We stress that EasyEdit2 is intended as a research tool for advancing understanding of model control, and must be used responsibly with proper safeguards.
@@ Acknowledgements
Our sincerest thanks are extended to CAA [^6] https://github.com/nrimsky/CAA, LM-Steer [^7] https://github.com/Glaciohound/LM-Steer, and AxBench [^8] https://github.com/stanfordnlp/axbench for their invaluable contributions to our project. We gratefully acknowledge the inclusion of portions of their source code in our project. We also extend our thanks to the community for its ongoing support and collaboration. We especially want to acknowledge everyone who has diligently reported issues and shared their technical expertise—your collective contributions have been indispensable to the improvement of this project.
This work was supported by the National Natural Science Foundation of China (No. 62576307), the Fundamental Research Funds for the Central Universities (226-2023-00138), Yongjiang Talent Introduction Programme (2021A-156-G), Ningbo Natural Science Foundation (2024J020), Tencent AI Lab Rhino-Bird Focused Research Program (RBFR2024003), and Information Technology Center and State Key Lab of CAD&CG.
@@ B.2 Adjustability of Steering Vectors
We investigate the adjustability of merged steering vectors on the Gemma-2-9b model by modulating their effect with different multipliers during inference. Using CAA-based vectors from safety and sentiment tasks, we test three merging strategies: Linear, TIES, and DARE-TIES. Details of these merging strategies are provided in Section 3.4. Each steering vector is scaled by a multiplier from 2-2 to $2$ and applied during inference.